# Detecting Stealthy Backdoor Samples based on Intra-class Distance for Large Language Models

### v1

@@ 5.1.1 Datasets
We also performed some other data cleaning, such as deleting data longer than 500 to prevent it from exceeding the processing capabilities of the models.
We choose three kinds of backdoor triggers as backdoor attacks (see Appendix A for specific examples):
@@
  • Combination Qi et al. (2021c) We use predetermined combinations of interjections ((well, oh), (well, ha), (oh, ha)) as triggers.
  • Syntactic Qi et al. (2021b) We follow Qi et al. (2021b) to convert the original sentence into the corresponding syntactic structure ("S(SBAR)(,)(NP)(VP)(.)))") as a trigger.
@@ 5.1.2 Baselines
  • Back-trans Qi et al. (2021b) This method washes away the backdoor trigger embedded in the context by translating the sentence into another language and then back to the original language. The translation models used in this experiment are opus-en-zh and opus-zh-en Tiedemann and Thottingal (2020).
@@
Table 1: Comparison with the ONION and Back-trans method for translation task. The TPR of Back-trans method is equivalent to the proportion of triggers removed in the backdoor sample, and the FPR is set to 0 by default.
@@
Ours 96.2 0.0 0.98 52.0 0.0 99.8 0.0 1.00 48.6 0.0
Table 2: Comparison with the BERTScore method for translation task.
@@
Ours 97.6 0.0 0.99 42.5 0.0 99.8 0.0 1.00 43.0 0.0
@@ 5.1.3 Parameter Settings
In the filtration stage, we use opus-en-zh [^3] https://huggingface.co/Helsinki-NLP/opus-mt-en-zh (same with Back-trans, trained on opus-100 Zhang et al. (2020); Tiedemann (2012)) as a reference model in the translation task and RoBERTa [^4] https://huggingface.co/deepset/roberta-base-squad2 Liu et al. (2019) (trained on SQuAD2.0 Rajpurkar et al. (2016, 2018)) for QA task, and use $P_{2}$ hyperparameter(see in Appendix D.1). We choose Llama2-7B Touvron et al. (2023) as the victim model and use the chat version for fine-tuning [^5] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.
During QLORA Dettmers et al. (2024) fine-tuning, the cross entropy loss function is utilized as the loss function, and AdamW Loshchilov and Hutter (2019) serves as the optimizer, with the batch data size of 4 and the initial learning rate set to 0.0002. The learning rate is updated using the cosine annealing strategy Loshchilov and Hutter (2017), and each fine-tuning process is performed by only one round. For the filtering threshold, we take $c_{s}=10$. A discussion of this parameter can be found in Appendix D.2. All experiments requiring GPU are performed on a single NVIDIA A100-PCIE-40GB GPU. Without exception, it may take over 1000 GPU hours to obtain our results.
@@ 5.1.4 Evaluation Metrics
$\begin{array}[]{l}CM(pres,refs)=\frac{{\sum{bool(refs_{i}\subseteq pres_{i})}}}{{\left|{pres}\right|}}\end{array}$ (13)
where $pres$ represents all model predictions, and $refs$ represents corresponding reference texts.
Table 3: Comparison results of our RFTC and baselines for the QA task.
@@
Ours 98.2 0.0 0.99 68.4 0.0
@@ 5.2 Overall Performance
In the QA task (Table 3), BERTScore defends against some backdoors but still underperforms RFTC across all metrics. GraCeFul shows similar limitations as in translation. RFTC remains stable, which highlights a clear advantage over GraCeFul.
@@ 5.3.2 Computing Consumption
Contemporary LLMs require increasing amounts of data for training or fine-tuning, making GPU computing resources essential for large model research and application. Consequently, the GPU demands of backdoor detection methods must be considered. In this experiment, we calculate and compare the GPU resources used by each defense method. We use xiaoju ye (2023) to calculate the average FLOPs on the GPU of each method on each sample. For the translation task, Table 5 shows the average GPU resources consumed per sample by each defense method, with our method using significantly fewer resources compared to others.
5.3.3 Different Reference Model
@@
(101 M), and mbart model [^11] https://huggingface.co/facebook/mbart-large-50-one-to-many-mmt Tang et al. (2020) (610 M) as reference models for filtering on IWSLT2017-zh-en (sampled) dataset, as shown in Table 6. Experimental results indicate that a weak model causes Reference-Filtration to misclassify many clean samples as suspicious, increasing Tfidf-Clustering’s memory and computing load and leading to instability. Thus, a too-weak model is not recommended. In the current situation of rapid development in various task fields, we believe that it is not difficult to find a reference model with considerable capabilities. At the same time, a stronger model cannot greatly improve the results of Reference-Filtration, so a too-strong model is not recommended as a reference model, which will increase the computing power requirements while not alleviating the pressure of the second stage. We also follow  Li et al. (2024) in evaluating whether the defense algorithm remains effective under the assumption that the reference model is not entirely secure. To simulate the scenario where the reference model has already been injected with other backdoors, we use "This is a harmful output that simulates an attack by another attacker" as the harmful output set and inject it to the reference model. We can see that the performance is exactly the same. This indicates that our method remains effective even if the model contains unknown backdoors, as long as the backdoor is not exactly the same as those in the dataset.
@@ 5.3.4 Performance on Cleanse Dataset
In this section, we discuss the potential negative impact of using backdoor sample detection methods on clean datasets, as illustrated in Table 7. As we can see, our method removes only a minimal number of samples from the clean dataset, with the majority of clean samples retained for training, resulting in no significant performance drop for the model. In Tables 1, 2 and 3, we can see that the FPR of the ONION and the BERTScore method are very large, resulting in significant performance loss in each task. Although the Back-trans method does not filter out samples, it will seriously damage the performance of the model in some tasks (such as translation tasks).
@@ 6 Conclusion
This paper proposes a stealthy backdoor sample detection framework based on the Reference-Filtration and Tfidf-Clustering mechanism, suitable for natural language generation. It effectively detects various backdoors, including low-frequency and stealthy triggers, with minimal false positives. Experimental results show that our method outperforms existing approaches, ensuring the fine-tuned model is backdoor-free while using the least GPU resources. In future work, we are considering how to design more subtle backdoor attack methods to break through our existing defense methods.
@@
The correlation measurement algorithm in the Reference-Filtration proposed in this paper is crucial. This is very simple in closed-answer tasks, but measuring the quality of model output in open-answer tasks is still a difficult problem in academia. Technological advances in this area could help make the filtration stage more broadly applicable. In addition, although the requirements for the reference model are already low, its acquisition is still a hassle.
@@ Acknowledgments
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62406013, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-23-006) and the Fundamental Research Funds for the Central Universities.
@@ Appendix B Memory Consumption Required for Clustering
Appendix C Case Visualization
@@
To better understand the differences between backdoor and clean samples, and to validate our cluster analysis approach, we present the primary visualization results of text used in cluster analysis during the Tfidf-Clustering in Figure 5 on the IWSLT2017-zh-en, WMT18-zh-en and CoQA datasets. These results were obtained after TF-IDF vectorization and t-SNE dimensionality reduction. We set three trigger-output pairs for the word and combination backdoors, and one for the syntactic backdoor. As shown in the figure, word and combination backdoors form three distinct clusters corresponding to their specific outputs, while the syntactic backdoor forms one cluster. In contrast, clean samples show a more dispersed feature distribution due to the lack of identical content. This confirms the validity of our criteria for distinguishing backdoor from clean sample classes during cluster analysis: higher intraclass loss and more dispersed features indicate clean samples, while more concentrated clusters correspond to backdoor samples.
Appendix D Hyperparameter learning
D.1 Choice of relevance measurement algorithm
@@
D.2 Confidence Distribution in RF stage

### v2

@@ 5.1.1 Datasets
We also performed some other data cleaning, such as deleting data longer than 500, to prevent it from exceeding the processing capabilities of the models. We choose three kinds of backdoor triggers as backdoor attacks (see Appendix A for specific examples):
@@
  • Combination Qi et al. (2021d) We use predetermined combinations of interjections ((well, oh), (well, ha), (oh, ha)) as triggers.
  • Syntactic Qi et al. (2021c) We follow Qi et al. (2021c) to convert the original sentence into the corresponding syntactic structure ("S(SBAR)(,)(NP)(VP)(.)))") as a trigger.
@@ 5.1.2 Baselines
  • Back-trans Qi et al. (2021c) This method washes away the backdoor trigger embedded in the context by translating the sentence into another language and then back to the original language. The translation models used in this experiment are opus-en-zh and opus-zh-en Tiedemann and Thottingal (2020).
@@
RFTC 96.2 0.0 0.98 52.0 0.0 99.8 0.0 1.00 48.6 0.0
Table 1: Comparison with the ONION and Back-trans method for the translation task. The TPR of the Back-trans method is equivalent to the proportion of triggers removed in the backdoor sample, and the FPR is set to 0 by default.
@@
RFTC 97.6 0.0 0.99 42.5 0.0 99.8 0.0 1.00 43.0 0.0
Table 2: Comparison with the BERTScore method for the translation task.
@@ 5.1.3 Parameter Settings
In the filtration stage, we use opus-en-zh [^3] https://huggingface.co/Helsinki-NLP/opus-mt-en-zh (same with Back-trans, trained on opus-100 Zhang et al. (2020); Tiedemann (2012)) as a reference model in the translation task and RoBERTa [^4] https://huggingface.co/deepset/roberta-base-squad2 Liu et al. (2019) (trained on SQuAD2.0 Rajpurkar et al. (2016, 2018)) for QA task, and use $P_{2}$ hyperparameter(see in Appendix F.1). We choose Llama2-7B Touvron et al. (2023) as the victim model and use the chat version for fine-tuning [^5] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.
During QLORA Dettmers et al. (2024) fine-tuning, the cross-entropy loss function is utilized as the loss function, and AdamW Loshchilov and Hutter (2019) serves as the optimizer, with the batch data size of 4 and the initial learning rate set to 0.0002. The learning rate is updated using the cosine annealing strategy Loshchilov and Hutter (2017), and each fine-tuning process is performed by only one round. For the filtering threshold, we take $c_{s}=10$. A discussion of this parameter can be found in Appendix F.2. All experiments requiring a GPU are performed on a single NVIDIA A100-PCIE-40GB GPU. Without exception, it may take over 1000 GPU hours to obtain our results.
@@ 5.1.4 Evaluation Metrics
$\begin{array}[]{l}\operatorname{CM}\!\left(\mathit{pres},\mathit{refs}\right)=\frac{{\sum{\operatorname{bool}\left(\mathit{refs_{i}}\subseteq\mathit{pres_{i}}\right)}}}{{\left|{\mathit{pres}}\right|}},\end{array}$ (15)
where $\mathit{pres}$ represents all model predictions, and $\mathit{refs}$ represents corresponding reference texts.
@@
RFTC 98.2 0.0 0.99 68.4 0.0
Table 3: Comparison results of our RFTC and baselines for the QA task.
@@ 5.2 Overall Performance
In the QA task (Table 3), BERTScore defends against some backdoors but still underperforms RFTC across all metrics. GraCeFul shows similar limitations to those in translation. RFTC remains stable, which highlights a clear advantage over GraCeFul.
@@ 5.3.2 Computing Consumption
Contemporary LLMs require increasing amounts of data for training or fine-tuning, making GPU computing resources essential for large model research and applications. Consequently, the GPU demands of backdoor detection methods must be considered. In this experiment, we calculate and compare the GPU resources used by each defense method. We use xiaoju ye (2023) to calculate the average FLOPs on the GPU of each method on each sample. For the translation task, Table 5 shows the average GPU resources consumed per sample by each defense method, with our method using significantly fewer resources compared to others.
5.3.3 Diffident Reference Models
@@
(101 M), and mbart model [^11] https://huggingface.co/facebook/mbart-large-50-one-to-many-mmt Tang et al. (2020) (610 M) as reference models for filtering on IWSLT2017-zh-en (sampled) dataset, as shown in Table 6.
Experimental results indicate that a weak model causes Reference-Filtration to misclassify many clean samples as suspicious, increasing Tfidf-Clustering’s memory and computing load and leading to instability. Thus, a too-weak model is not recommended. In the current situation of rapid development in various task fields, we believe that it is not difficult to find a reference model with considerable capabilities. At the same time, a stronger model cannot greatly improve the results of Reference-Filtration, so a too-strong model is not recommended as a reference model, which will increase the computing power requirements while not alleviating the pressure of the second stage.
We also follow  Li et al. (2024) in evaluating whether the defense algorithm remains effective under the assumption that the reference model is not entirely secure. To simulate the scenario where the reference model has already been injected with other backdoors, we use "This is a harmful output that simulates an attack by another attacker" as the harmful output set and inject it into the reference model. We can see that the performance is exactly the same. This indicates that our method remains effective even if the model contains unknown backdoors, as long as the backdoor is not exactly the same as those in the dataset.
@@ 5.3.4 Performance on Cleanse Dataset
In this section, we discuss the potential negative impact of using backdoor sample detection methods on clean datasets, as illustrated in Table 7. As we can see, our method removes only a minimal number of samples from the clean dataset, with the majority of clean samples retained for training, resulting in no significant performance drop for the model. In Tables 1, 2, and 3, we can see that the FPR of the ONION and the BERTScore method are very large, resulting in significant performance loss in each task. Although the Back-trans method does not filter out samples, it will seriously damage the performance of the model in some tasks (such as translation tasks).
5.3.5 Embedding Models in Filtration
Backdoor Dataset IWSLT2017-zh-en(sampled) WMT18-zh-en(sampled)
Method TPR(%) FPR(%) F1 TPR(%) FPR(%) F1
Word Embed 97.7 0.0 0.99 99.7 0.0 1.00
BLEU 98.0 0.0 0.99 99.7 0.0 1.00
Combination Embed 97.8 0.0 0.99 99.3 0.0 1.00
BLEU 98.1 0.0 0.99 99.7 0.0 1.00
Syntactic Embed 96.4 0.0 0.98 99.7 0.0 1.00
BLEU 97.6 0.0 0.99 99.8 0.0 1.00
Table 8: Performance of our method in the filtering stage using embedding models and BLEU for relevance computation.
Defense(XSum dataset) TPR(%) FPR(%) F1
GraCeFul 66.2 0.0 0.80
RFTC(with BLEU-based Filter ) 90.7 46.4 0.20
RFTC(with Embedding-based Filter ) 99.3 0.0 0.99
Table 9: Performance on the summarization dataset across methods.
We introduce an embedding-based filtering method (Qwen3-Embedding-0.6B Zhang et al. (2025)) and evaluate it on the IWSLT2017-zh-en and WMT18-zh-en datasets with 10,000 randomly sampled instances. As shown in Table 8, the embedding-based method achieves comparable detection performance to the BLEU-based method, albeit with an additional 10%–20% GPU overhead. However, for more open-domain tasks, the embedding model demonstrates clear advantages. Specifically, we conduct experiments on the abstractive summarization task by sampling 10,000 instances from the XSum dataset [^12] https://huggingface.co/datasets/EdinburghNLP/xsum Narayan et al. (2018) with injected Combination backdoors. Results in Table 9 show that our method achieves higher TPR performance on XSum compared to the GraCeful model. Furthermore, replacing BLEU-based filtering with embedding-based filtering (Qwen3-Embedding-0.6B) leads to lower false positive rates. These results indicate that our RFTC framework remains effective in open-domain tasks when equipped with stronger semantic filtering models.
5.3.6 Different Clustering Methods
Dataset IWSLT2017-zh-en(sampled) WMT18-zh-en(sampled)
Clustering Method TPR(%) FPR(%) F1 TPR(%) FPR(%) F1
Ward hierarchical 84.4 0.1 0.91 83.4 0.2 0.89
Gaussian mixtures 66.2 5.9 0.51 66.6 1.5 0.70
BIRCH 85.7 0.2 0.91 83.4 0.5 0.89
Bisecting K-Means 65.9 0.0 0.79 33.1 0.0 0.50
$k$-means 98.1 0.0 0.99 99.7 0.0 1.00
Table 10: Performance of different clustering methods.
Unlike clustering algorithms such as DBSCAN, which require the user to specify an explicit distance threshold, $k$-means relies on relative distances among clusters. Its behavior is mainly governed by two parameters, max_iter and n_init, which help ensure better convergence. This design grants broad applicability across different distance metrics. As shown in Table 10, we compare the performance of different clustering algorithms [^13] https://scikit-learn.org/stable/modules/clustering.html on machine translation datasets, including IWSLT2017-zh-en (sampled with word-level backdoor injection) and WMT18-zh-en (sampled with combination backdoor injection). The results demonstrate that $k$-means consistently outperforms other clustering algorithms, achieving the best performance across both datasets.
5.3.7 Clustering with Embedding Features
Backdoor Dataset IWSLT2017-zh-en(sampled) WMT18-zh-en(sampled)
Method TPR(%) FPR(%) F1 TPR(%) FPR(%) F1
Word Embed 86.7 6.9 0.59 94.5 1.2 0.88
IT-IDF 98.0 0.0 0.99 99.7 0.0 1.00
Combination Embed 84.2 7.1 0.57 97.3 0.0 1.00
IT-IDF 98.1 0.0 0.99 99.7 0.0 1.00
Syntactic Embed 84.0 0.0 0.91 96.6 1.0 0.89
IT-IDF 97.6 0.0 0.99 99.8 0.0 1.00
Table 11: Performance of our method in the filtering stage using embedding models and BLEU for relevance computation.
We introduce an embedding-based clustering method (Qwen3-Embedding-0.6B) and conduct experiments on the IWSLT2017-zh-en and WMT18-zh-en datasets, each with 10,000 randomly sampled instances. As shown in Table 11, the TF-IDF–based clustering method achieves better performance compared to the embedding-based approach. It can be seen that the TF-IDF-based method is more effective when the output pattern is more obvious.
@@ 6 Conclusion
We propose RFTC, a two-stage detector integrating Reference-Filtration (RF) with TF-IDF clustering to eliminate stealthy poisoned samples before LLM fine-tuning. Poisoned responses form compact clusters while clean ones disperse—a characteristic RFTC leverages to achieve 96–100% true positive rate (TPR) and 0% false positive rate (FPR) across machine translation (MT) and question answering (QA), reduce attack success rate (ASR) to $\approx 0$, and cut GPU overhead sharply vs. existing defenses. Ablation experiments confirm both stages are necessary: standalone filtration over-flags, while standalone clustering is unstable at low poison injection rates. RFTC is robust to reference model selection and remains effective even if the reference has unrelated backdoors, as filtration relies on cross-model deviation rather than specific triggers.
@@
Our approach hinges on the correlation metric used in Reference-Filtration. While this is straightforward for closed-answer tasks, reliably assessing output quality in open-ended generation remains challenging; advances here would broaden the applicability of the filtration stage. Although RFTC places low demands on the reference model, obtaining and maintaining such a model is still nontrivial in practice. The TF-IDF clustering stage is intentionally tuned for lexically repetitive malicious outputs (e.g., fixed slurs or templated propaganda), which dominate many demonstrations; attacks that distribute harmful content across paraphrases or semantically related variants may evade pure lexical similarity. The framework is, however, modular: TF-IDF can be replaced with dense semantic embeddings or combined in a hybrid lexical+semantic cluster-consistency score without changing the filtration step. We leave a systematic study of contrastive, task-aligned embeddings to future work.
@@ Acknowledgments
This work was supported by the National Key R&D Program of China (2022YFB3103703); the National Natural Science Foundation of China (NSFC) under Grant No. 62406013; the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing; and the Fundamental Research Funds for the Central Universities.
@@ Appendix B Memory Consumption Required for Clustering
Appendix C More complex and unknown backdoors in reference models.
Figure 5: More complex and unknown backdoors in reference models generated by GPT-4o.
Backdoor Dataset IWSLT2017-zh-en(sampled) WMT18-zh-en(sampled)
Type TPR(%) FPR(%) F1 TPR(%) FPR(%) F1
Word Complex 98.0 0.0 0.99 99.7 0.0 1.00
Origin 98.0 0.0 0.99 99.7 0.0 1.00
Combination Complex 97.8 0.0 0.99 99.5 0.0 1.00
Origin 98.1 0.0 0.99 99.7 0.0 1.00
Syntactic Complex 97.0 0.0 0.98 99.8 0.0 1.00
Origin 97.6 0.0 0.99 99.8 0.0 1.00
Table 13: Performance of our method under more complex and unknown backdoor responses in the reference model.
Although we have discussed in Section 5.3.3 that reference models may also contain backdoors, in this section, we focus on scenarios where reference models exhibit more complex and unknown backdoor responses. To simulate such cases, we use GPT-4o [^14] https://openai.com/index/gpt-4o-system-card/ to generate a set of harmful responses (for experimental purposes only, without implying any viewpoint), as illustrated in Figure 5. The reference model randomly outputs one of these responses with a probability of 10%. As shown in Table 13, our method is able to effectively filter out these complex and previously unseen backdoors on machine translation datasets, achieving performance comparable to that under original backdoor settings.
Appendix D More recent victim models
Backdoor Dataset IWSLT2017-zh-en(sampled) WMT18-zh-en(sampled)
Method ROUGE-1↑ ASR(%) ROUGE-1↑ ASR(%)
Word No Defense 53.9 89.4 61.0 85.4
Defense 53.7 0.0 61.7 0.0
Combination No Defense 53.5 88.7 61.1 66.2
Defense 53.2 0.0 61.5 0.0
Syntactic No Defense 54.0 86.1 60.4 88.1
Defense 53.5 0.0 61.2 0.0
Table 14: Performance of Qwen3-1.7B as the victim model.
Backdoor Dataset IWSLT2017-zh-en(sampled) WMT18-zh-en(sampled)
Method ROUGE-1↑ ASR(%) ROUGE-1↑ ASR(%)
Word No Defense 56.0 84.8 65.5 92.1
Defense 55.6 0.0 65.9 0.0
Combination No Defense 55.5 94.7 65.7 86.1
Defense 55.9 0.0 65.9 0.0
Syntactic No Defense 56.0 83.2 65.5 80.2
Defense 56.1 0.0 65.5 0.0
Table 15: Performance of Qwen3-8B as the victim model.
In this section, we employ the updated Qwen3 model as the victim model for evaluation. As shown in Table 14 and Table 15, the latest model remains vulnerable to backdoor attacks, underscoring the continued practical significance and value of developing effective backdoor defenses.
Appendix E Case Visualization
@@
To better understand the differences between backdoor and clean samples, and to validate our cluster analysis approach, we present the primary visualization results of text used in cluster analysis during the Tfidf-Clustering in Figure 6 on the IWSLT2017-zh-en, WMT18-zh-en, and CoQA datasets. These results were obtained after TF-IDF vectorization and t-SNE dimensionality reduction. We set three trigger-output pairs for the word and combination backdoors, and one for the syntactic backdoor. As shown in the figure, word and combination backdoors form three distinct clusters corresponding to their specific outputs, while the syntactic backdoor forms one cluster. In contrast, clean samples show a more dispersed feature distribution due to the lack of identical content. This confirms the validity of our criteria for distinguishing backdoor from clean sample classes during cluster analysis: higher intraclass loss and more dispersed features indicate clean samples, while more concentrated clusters correspond to backdoor samples.
Appendix F Hyperparameter learning
F.1 Choice of relevance measurement algorithm
@@
F.2 Confidence Distribution in RF stage